Steven had a balance of $1,200 in his first stare bank checking account. He deposits a $387.89 paycheck, a $437.12 dividend check, and a personal check for $250 into his account. He wants to receive $400 in cash. How much will he have in his account after the transaction?


1
Expert's answer
2021-09-20T23:28:09-0400

Steven had $1,200 in his checking account. Then he deposits a $387.89

Hence have "1200+387.89=1587.80 dollars"

Also he puts j$437.12 dividend check, and a personal check for $250

Therefore his account now has


"1587.80+437.12+250=2275.01"


But because he wants to obtain cash $400, he will have

"2275.01-400=1875.01dollars"


Hence after all those transactions Steven will have on his account "1875.01dollars"